OpenStudy (anonymous):

Medals! a^2 + 12ab + 27b^2

OpenStudy (anonymous):

What are you wanting to do with it?

OpenStudy (kainui):

Perhaps it's an alchemist's way of saying what proportion of metals are involved in an alloy, I can only speculate upon this vague description, for a philosopher I am not.

OpenStudy (anonymous):

I want to find the answer

OpenStudy (anonymous):

It says factor it, but I have trouble doing these

OpenStudy (anonymous):

do you know how to solve this?

OpenStudy (kainui):

Yeah so the number in the middle is two numbers added together and the number on the right is two numbers multiplied. (a+xb)(a+yb)=a^2+(x+y)ab+(xy)b^2 So now you can see x+y=12 and xy=27. You have two equations and two unknowns, you can solve this by plugging one in the other.

OpenStudy (anonymous):

ok, I'm a bit lost, are you sure this is the right way, this method I haven't seen before, I've only seen the Quadratic formula?